OpenStudy (anonymous):

The local day care facility has a ratio of 1 adult for every 6 children. How man adults are working if there are 42 children?

OpenStudy (anonymous):

7 adults

OpenStudy (anonymous):

1/6 = x/42

OpenStudy (anonymous):

\[\frac{1}{6} = \frac{x}{42}\]

OpenStudy (anonymous):

x = 7

OpenStudy (mathteacher1729):

1 adult = 6 children 2 adults = 12 children 3 adults = 18 children 4 adults = 24 children ... Note how I'm just multiplying the number adults by 6 to find the number of children. OR Note how I'm just DIVIDING the number of children by 6 to find the number of adults. :)